# reference.bib

### v1

@@ III-B Cross-modal Feature Restoration
Existing methods for event-image fusion in low-light scenarios often fail to balance computational efficiency and robust cross-modal interaction, particularly in capturing high-frequency motion cues from events and preserving spatial details from images [jiang2023event, liang2024towards]. Furthermore, they frequently treat event and image data as independent streams, failing to fully exploit the fine-grained spatiotemporal relationships between them. To address this, we propose the Cross-RWKV Module, which integrates Re-WKV Attention for capturing global interactions and Cross-Modal Switch Omni-Shift (CS-Shift) for local feature enhancement and multi-modal information interaction. Specifically, multiple Cross-RWKV Modules form a 4-level encoder-decoder. The encoder progressively downsamples spatial dimensions and expands channels, while the decoder upsamples to restore resolution, enabling multi-scale cross-modal fusion. The outputs $O_{\text{img}},O_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2C}$ is split by a 3×3 convolution into $\hat{X}_{\text{img}},\hat{X}_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$. As shown in Fig.3, the module operates through spatial and channel mix, enabling efficient cross-modal fusion of event and image features. The details of the Cross-RWKV Block are as follows:
@@ III-B1 Spatial Mix
Following the Restore-RWKV [yang2024restore], we employ the Re-WKV mechanism to compute the global attention result $\text{wkv}\in\mathbb{R}^{T\times C}$. At the same time, key representations are directly exchanged between image tokens and event tokens, allowing each modality to attend to the spatial structure of the other modality, thereby enabling joint spatial reasoning between images and events:
@@ III-C Frequency-aware Feature Fusion
Given the restored image feature $\hat{X}_{\text{img}}$, event feature $\hat{X}_{\text{ev}}$, and raw image input $X_{\text{inp}}$, we first concatenate them along the channel dimension to form a unified representation:
$X_{\text{fused}}=\text{Concat}(\hat{X}_{\text{img}},\hat{X}_{\text{ev}},X_{\text{inp}})\in\mathbb{R}^{3C\times H\times W}.$
@@
$G_{\sigma_{c}}(x,y)=\frac{1}{2\pi\sigma_{c}^{2}}\exp\left(-\frac{x^{2}+y^{2}}{2\sigma_{c}^{2}}\right),$
@@
$\hat{X}_{\text{freq},c}=\mathcal{F}^{-1}\left(\mathcal{F}(X_{\text{freq},c})\cdot\mathcal{F}(G_{\sigma_{c}})\right),$
@@
$\hat{X}_{\text{spat}}=\text{DeformConv}(X_{\text{spat}}),$
@@ III-D Loss Function
We optimise EvRWKV in an end‑to‑end fashion by minimising a weighted combination of four complementary losses: the pixel‑wise reconstruction loss $\mathcal{L}_{r}$, the perceptual loss $\mathcal{L}_{p}$, the Structural Similarity loss $\mathcal{L}_{s}$, and its multi‑scale extension $\mathcal{L}_{m}$. Specifically,
@@
$\mathcal{L}_{r}=\sqrt{\bigl{\lVert}I_{\mathrm{en}}-I_{\mathrm{gt}}\bigr{\rVert}+\varepsilon^{2}},$ (22)
@@
We employ the perceptual loss [ledig2017photo] to preserve the high-level similarity between the fused image and source images, which is given by
@@
where $\Phi$ represents feature extraction using the Alex network [krizhevsky2012imagenet].
For every $11\times 1$ 11 Gaussian window $\Omega$ we compute the Structural Similarity (SSIM) index between the enhanced patch $x=I_{\mathrm{en}}\!\big{|}_{\Omega}$ and the corresponding ground-truth patch $y=I_{\mathrm{gt}}\!\big{|}_{\Omega}$:
@@ IV EXPERIMENTS
Method Publication Backbone Inputs SDE-in SDE-out RELED
[…]
SNR-Net[xu2022snr] CVPR2022 Transformer $\checkmark$ $\times$ 20.05 0.630 22.18 0.661 26.47 0.851
Uformer[wang2022uformer] CVPR2022 Transformer $\checkmark$ $\times$ 21.09 0.752 22.32 0.725 28.86 0.832
LLFlow-L-SKF[wu2023learning] CVPR2023 CNN $\checkmark$ $\times$ 20.92 0.661 21.68 0.647 28.62 0.862
Retinexformer[cai2023retinexformer] ICCV2023 Transformer $\checkmark$ $\times$ 21.3 0.692 22.92 0.683 26.66 0.865
eSL-Net[wang2020event] ECCV2020 CNN $\checkmark$ $\checkmark$ 21.42 0.725 21.39 0.681 25.76 0.796
ELIE[jiang2023event] TMM2023 Transformer $\checkmark$ $\checkmark$ 19.98 0.617 20.69 0.653 26.62 0.862
LLVE-SEG[liu2023low] AAAI2023 Transformer $\checkmark$ $\checkmark$ 21.79 0.705 22.35 0.690 29.19 0.875
ELEDNet[kim2024towards] ECCV2024 Transformer $\checkmark$ $\checkmark$ 21.46 0.7126 22.91 0.726 31.96 0.910
EvLight[liang2024towards] CVPR2024 Transformer $\checkmark$ $\checkmark$ 22.21 0.758 22.13 0.725 31.29 0.865
Ours - RWKV $\checkmark$ $\checkmark$ 23.09 0.766 23.43 0.767 32.18 0.910
@@ IV-A1 Datasets and Evaluation Metrics
SED [liang2024towards], SDSD [wang2021seeing], and RELED [kim2024towards] datasets are selected to test the enhancement performance of our EvRWKV framework.
SED dataset contains 91 image–event paired sequences (43 indoor and 48 outdoor) captured with a DAVIS346 event camera at a resolution of $346\times 260$; it is primarily used to evaluate the capability of different methods to leverage event cues for low-light enhancement under diverse illumination conditions.
SDSD dataset provides 150 low-light/normal-light paired video sequences with an original resolution of $1920\times 1080$; following the dynamic-subset split (125 sequences for training and 25 for testing), all videos are down-sampled to $346\times 260$, and their event streams are synthesized with the v2e [hu2021v2e] simulator. This dataset evaluates the robustness of enhancement methods in dynamic low-light scenes.
@@
Together, these three benchmarks provide a comprehensive testbed for assessing the multi-modal low-light enhancement capability of our EvRWKV method. We use the peak-signal-to-noise ratio (PSNR) [hore2010image] and SSIM [wang2004image] for evaluation.
@@
[…]
SNR-Net[xu2022snr] CVPR’22 24.74 0.830 24.82 0.740
Uformer[wang2022uformer] CVPR’22 24.03 0.900 24.08 0.818
LLFlow-L-SKF[wu2023learning] CVPR’23 23.39 0.818 20.39 0.634
Retinexformer[cai2023retinexformer] ICCV’23 25.90 0.852 26.08 0.815
@@
ELIE[jiang2023event] TMM’23 27.46 0.879 23.29 0.742
eSL-Net[wang2020event] ECCV’20 23.68 0.821 24.95 0.805
LLVE-SEG[liu2023low] AAAI’23 27.58 0.888 23.51 0.726
ELEDNet[kim2024towards] ECCV’24 28.52 0.912 25.08 0.828
EvLight[liang2024towards] CVPR’24 28.49 0.909 25.17 0.817
@@ IV-A2 Implementation Details
All experiments were conducted with the Adam optimizer [kingma2014adam]. Learning rates were set to $1e{-4}$ for SDE datasets, $1.5e{-4}$ for SDSD datasets, and $1e{-4}$ for RELED datasets. The framework was trained on an NVIDIA RTX 3090 GPU for 80 training cycles with a batch size of eight. For the training set, we employed data-augmentation techniques consisting of random $256\times 256$ crops, horizontal flips, and rotations of $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. For the test set, we applied only a center crop of $256\times 256$.
@@ IV-A3 State-of-the-Art Methods for Comparisons
We compared our EvRWKV with ten methods, including five methods that only use RGB images as input(SNR-Net [xu2022snr], Uformer [wang2022uformer], LLFlow-L-SKF [wu2023learning], RetinexMac [10810469], Retinexformer [cai2023retinexformer]), five methods that use a RGB image and paired events as inputs(ELIE [jiang2023event], eSL-Net [wang2020event], LLVE-SEG [liu2023low], ELEDNet [kim2024towards], EvLight [liang2024towards]). Since the source code of LLVE-SEG [liu2023low] is not publicly available, we implemented it according to the description in its paper. For the SNR-Net [xu2022snr], Uformer [wang2022uformer], LLFlow-L-SKF [wu2023learning], RetinexMac [10810469], Retinexformer [cai2023retinexformer], ELIE [jiang2023event], eSL-Net [wang2020event], ELEDNet [kim2024towards] and EvLight [liang2024towards] methonds, we used the codes released by the authors to output their results.
@@ IV-B Quantitative and Qualitative Evaluation
We evaluate the proposed EvRWKV framework on three challenging datasets—SDE, SDSD, and RELED—covering diverse low-light enhancement scenarios. SDE and RELED contain real-world event-image pairs captured under indoor/outdoor and low-light blurry conditions respectively, while SDSD consists of synthetic events generated from dynamic video sequences. We use PSNR and SSIM as quantitative metrics and provide detailed qualitative analysis to validate the effectiveness of our approach.
@@ IV-B1 Evaluation on the SDE Dataset
For low-light image enhancement on the SDE dataset, the ability to simultaneously suppress noise and preserve fine structural details is the key evaluation criterion. We first qualitatively compare our EvRWKV method with several representative approaches including eSL-Net [wang2020event], ELEDNet [kim2024towards], EvLight [liang2024towards], and RetinexMac [10810469]. As shown in Fig. 5, RetinexMac and ELEDNet reduce noise to some extent but tend to oversmooth textures, resulting in loss of important details in dark regions. As shown in Fig. 5, EvLight maintains more structural information but suffers from residual noise and blur artifacts. eSL-Net, with a relatively small model size, provides limited enhancement capabilities and often underperforms in challenging scenes. In contrast, our EvRWKV method achieves a better balance between noise suppression and texture preservation, producing cleaner, brighter images with rich details and minimal artifacts. This demonstrates the effectiveness of our dual-domain fusion and adaptive gating mechanisms in handling complex low-light scenarios.
In terms of quantitative evaluation, Table I lists the PSNR and SSIM scores of different methods on the SDE dataset under both SDE-in and SDE-out settings. Our method achieves the highest scores of 23.09 PSNR and 0.766 SSIM for SDE-in, as well as 23.43 PSNR and 0.767 SSIM for SDE-out. These results outperform all compared methods including SNR-Net [xu2022snr], Uformer [wang2022uformer], LLFlow-L-SKF [wu2023learning], LLVE-SEG [liu2023low], Retinexformer [cai2023retinexformer], ELIE [jiang2023event], LLVE-SEG [liu2023low], and the aforementioned qualitative comparison models. The significant improvements in both perceptual quality and objective metrics validate the superiority of our EvRWKV framework in low-light image enhancement.
@@ IV-B2 Evaluation on the SDSD Dataset
We further qualitatively compare the performance of different methods on the SDSD dataset for low-light image enhancement under challenging lighting and motion conditions. In Fig. 6, we observe that most methods improve the visibility and brightness of degraded images. However, RetinexMac [10810469], ELEDNet [kim2024towards], and EvLight [liang2024towards] exhibit limitations: As shown in Fig. 6, RetinexMac and ELEDNet tend to over-smooth textures and lose fine details, While EvLight, although preserving more structural information, causes noticeable color distortion in some regions. As shown in Fig. 6, eSL-Net [wang2020event] shows relatively limited enhancement capability due to its smaller model size. In comparison, our EvRWKV method delivers superior noise suppression and detail preservation, effectively enhancing both static and dynamic scene elements without introducing artifacts. The enhanced images show clearer textures, better contrast, and natural brightness, demonstrating the robustness of our dual-domain cross-modal fusion strategy.
@@ IV-B3 Evaluation on the RELED Dataset
To comprehensively evaluate the enhancement performance of our EvRWKV method on the RELED dataset, we compare it qualitatively with several representative approaches including eSL-Net [wang2020event], ELEDNet [kim2024towards], EvLight [liang2024towards], and RetinexMac [10810469]. As shown in Fig. 7, the RELED dataset contains extremely low-light images with severe motion blur and noise. As shown in Fig. 7, eSL-Net often causes overexposure while attempting to enhance brightness. RetinexMac boosts overall brightness but fails to eliminate noticeable blur. ELEDNet and EvLight aim to balance illumination, but this often leads to the loss of finer details. In contrast, our EvRWKV method effectively leverages event-based motion cues and image textures to suppress noise, reduce blur, and restore fine details, producing visually clearer, sharper, and more natural images without color artifacts or overexposure.
Table I reports the PSNR and SSIM scores for different methods on the RELED dataset. Our EvRWKV achieves the best quantitative results with a PSNR of 32.18 and SSIM of 0.910, significantly outperforming other methods including eSL-Net, ELEDNet, EvLight, and RetinexMac. These quantitative gains demonstrate the robustness and generalization capability of our dual-domain cross-modal fusion framework in handling challenging low-light and motion-degraded scenarios. Overall, our method achieves excellent qualitative and quantitative results across all three real-world low-light enhancement benchmarks.
Figure 8: Visualization of ablation results.
IV-C Ablation Study
We conduct comprehensive ablation studies on the SDSD-in dataset to evaluate the contribution of each key component in our EvRWKV framework. In particular, we assess the effectiveness of the EISFE module, the SpatialMix and ChannelMix mechanisms within the Cross-RWKV module, and the role of each individual loss term in the overall objective function. For each study, we remove or modify a single module while keeping the rest of the framework unchanged to isolate its impact. The results clearly demonstrate the necessity and complementarity of these components in achieving robust and high-quality low-light enhancement. As shown in Fig. 8, the removal of EISFE (b) results in color distortion. Without SpatialMix (c), visible noise and blurred textures appear, while omitting ChannelMix (d) leads to flat tonal rendering. The full model (e) demonstrates the best visual clarity and color fidelity, especially in complex shadow regions.
IV-C1 Impact of EISFE Module
IV-C2 Impact of SpatialMix and ChannelMix

### v2

@@ III-B Cross-modal Feature Restoration
Existing methods for event-image fusion in low-light scenarios often fail to balance computational efficiency and robust cross-modal interaction, particularly in capturing high-frequency motion cues from events and preserving spatial details from images [17, 30]. Furthermore, they frequently treat event and image data as independent streams, failing to fully exploit the fine-grained spatiotemporal relationships between them. To address this, we propose the Cross-RWKV Module, which integrates Re-WKV Attention for capturing global interactions and Cross-Modal Switch Omni-Shift (CS-Shift) for local feature enhancement and multi-modal information interaction. Specifically, multiple Cross-RWKV Modules form a 4-level encoder-decoder. The encoder progressively downsamples spatial dimensions and expands channels, while the decoder upsamples to restore resolution, enabling multi-scale cross-modal fusion. The outputs $O_{\text{img}},O_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times 2C}$ is split by a 3×3 convolution into $\hat{X}_{\text{img}},\hat{X}_{\text{ev}}\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$. As shown in Fig.4, the module operates through spatial and channel mix, enabling efficient cross-modal fusion of event and image features. The details of the Cross-RWKV Block are as follows:
@@ III-B1 Spatial Mix
Following the Restore-RWKV [57], we employ the Re-WKV mechanism to compute the global attention result $\text{wkv}\in\mathbb{R}^{T\times C}$. At the same time, key representations are directly exchanged between image tokens and event tokens, allowing each modality to attend to the spatial structure of the other modality, thereby enabling joint spatial reasoning between images and events:
@@ III-C Frequency-aware Feature Fusion
Method Publication Backbone Inputs SDE-in SDE-out RELED
[…]
SNR-Net[55] CVPR2022 Transformer $\checkmark$ $\times$ 20.05 0.630 22.18 0.661 26.47 0.851
Uformer[51] CVPR2022 Transformer $\checkmark$ $\times$ 21.09 0.752 22.32 0.725 28.86 0.832
LLFlow-L-SKF[53] CVPR2023 CNN $\checkmark$ $\times$ 20.92 0.661 21.68 0.647 28.62 0.862
Retinexformer[3] ICCV2023 Transformer $\checkmark$ $\times$ 21.3 0.692 22.92 0.683 26.66 0.865
eSL-Net[46] ECCV2020 CNN $\checkmark$ $\checkmark$ 21.42 0.725 21.39 0.681 25.76 0.796
ELIE[17] TMM2023 Transformer $\checkmark$ $\checkmark$ 19.98 0.617 20.69 0.653 26.62 0.862
LLVE-SEG[35] AAAI2023 Transformer $\checkmark$ $\checkmark$ 21.79 0.705 22.35 0.690 29.19 0.875
ELEDNet[20] ECCV2024 Transformer $\checkmark$ $\checkmark$ 21.46 0.7126 22.91 0.726 31.96 0.910
EvLight[30] CVPR2024 Transformer $\checkmark$ $\checkmark$ 22.21 0.758 22.13 0.725 31.29 0.865
Ours - RWKV $\checkmark$ $\checkmark$ 23.09 0.766 23.43 0.767 32.18 0.910
@@
Given the restored image feature $\hat{X}_{\text{img}}$, event feature $\hat{X}_{\text{ev}}$, and raw image input $X^{\text{img}}$, we first concatenate them along the channel dimension to form a unified representation:
$X_{\text{fused}}=\text{Concat}(\hat{X}_{\text{img}},\hat{X}_{\text{ev}},X^{\text{img}}).$ (19)
@@
$G_{\sigma_{c}}(x,y)=\frac{1}{2\pi\sigma_{c}^{2}}\exp\left(-\frac{x^{2}+y^{2}}{2\sigma_{c}^{2}}\right),$ (20)
@@
$\hat{X}_{\text{freq},c}=\mathcal{F}^{-1}\left(\mathcal{F}(X_{\text{freq},c})\cdot\mathcal{F}(G_{\sigma_{c}})\right),$ (21)
@@
$\hat{X}_{\text{spat}}=\text{DeformConv}(X_{\text{spat}}),$ (22)
@@ III-D Loss Function
We optimise EvRWKV in an end‑to‑end fashion by minimising a weighted combination of four complementary losses: the pixel‑wise reconstruction loss $\mathcal{L}_{r}$, the perceptual loss $\mathcal{L}_{p}$, the Structural Similarity loss $\mathcal{L}_{s}$, and its multi‑scale extension $\mathcal{L}_{m}$. Specifically, $\mathcal{L}_{\text{total}}$ is given by
@@
$\mathcal{L}_{r}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(I_{\text{en}}-I_{\text{gt}})_{i}^{2}+\epsilon^{2}},$ (26)
@@
We employ the perceptual loss [25] to preserve the high-level similarity between the fused image and source images, which is given by
@@
where $\Phi$ represents feature extraction using the Alex network [22].
For every $11\times 11$ Gaussian window $\Omega$ we compute the Structural Similarity (SSIM) index between the enhanced patch $x=I_{\mathrm{en}}\!\big{|}_{\Omega}$ and the corresponding ground-truth patch $y=I_{\mathrm{gt}}\!\big{|}_{\Omega}$:
@@ IV-A1 Datasets and Evaluation Metrics
SED [30], SDSD [47], and RELED [20] datasets are selected to test the enhancement performance of our EvRWKV framework.
SED dataset contains 91 image–event paired sequences (43 indoor and 48 outdoor) captured with a DAVIS346 event camera at a resolution of $346\times 260$; it is primarily used to evaluate the capability of different methods to leverage event cues for LLIE under diverse illumination conditions.
SDSD dataset provides 150 low-light/normal-light paired video sequences with an original resolution of $1920\times 1080$; following the dynamic-subset split (125 sequences for training and 25 for testing), all videos are down-sampled to $346\times 260$, and their event streams are synthesized with the v2e [15] simulator. This dataset evaluates the robustness of enhancement methods in dynamic low-light scenes.
@@
Together, these three benchmarks provide a comprehensive testbed for assessing the multi-modal LLIE capability of our EvRWKV method. We use the peak-signal-to-noise ratio (PSNR) [14] and SSIM [52] for evaluation.
@@
$uparrow PSNR\uparrow SSIM\uparrow$
SNR-Net[55] CVPR’22 24.74 0.830 24.82 0.740
Uformer[51] CVPR’22 24.03 0.900 24.08 0.818
LLFlow-L-SKF[53] CVPR’23 23.39 0.818 20.39 0.634
Retinexformer[3] ICCV’23 25.90 0.852 26.08 0.815
@@
ELIE[17] TMM’23 27.46 0.879 23.29 0.742
eSL-Net[46] ECCV’20 23.68 0.821 24.95 0.805
LLVE-SEG[35] AAAI’23 27.58 0.888 23.51 0.726
ELEDNet[20] ECCV’24 28.52 0.912 25.08 0.828
EvLight[30] CVPR’24 28.49 0.909 25.17 0.817
@@ IV-A2 Implementation Details
All experiments were conducted with the Adam optimizer [21]. Learning rates were set to 1e41e{-4} for SDE datasets, 1.5e41.5e{-4} for SDSD datasets, and 1e41e{-4} for RELED datasets. The framework was trained on an NVIDIA RTX 3090 GPU for 80 training cycles with a batch size of eight. For the training set, we employed data-augmentation techniques consisting of random $256\times 256$ crops, horizontal flips, and rotations of $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. For the test set, we applied only a center crop of $256\times 256$.
@@ IV-A3 State-of-the-Art Methods for Comparisons
We compared our EvRWKV with ten methods, including five methods that only use RGB images as input(SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], RetinexMac [33], Retinexformer [3]), five methods that use a RGB image and paired events as inputs(ELIE [17], eSL-Net [46], LLVE-SEG [35], ELEDNet [20], EvLight [30]). Since the source code of LLVE-SEG [35] is not publicly available, we implemented it according to the description in its paper. For the SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], RetinexMac [33], Retinexformer [3], ELIE [17], eSL-Net [46], ELEDNet [20] and EvLight [30] methonds, we used the codes released by the authors to output their results.
@@ IV-B Quantitative and Qualitative Evaluation
We evaluate the proposed EvRWKV framework on three challenging datasets—SDE, SDSD, and RELED—covering diverse LLIE scenarios. SDE and RELED contain real-world event-image pairs captured under indoor/outdoor and low-light blurry conditions respectively, while SDSD consists of synthetic events generated from dynamic video sequences. We use PSNR and SSIM as quantitative metrics and provide detailed qualitative analysis to validate the effectiveness of our approach.
@@ IV-B1 Evaluation on the SDE Dataset
For LLIE on the SDE dataset, the ability to simultaneously suppress noise and preserve fine structural details is the key evaluation criterion. We first qualitatively compare our EvRWKV method with several representative approaches including eSL-Net [46], ELEDNet [20], EvLight [30], and RetinexMac [33]. As shown in Fig. 6, RetinexMac and ELEDNet reduce noise to some extent but tend to oversmooth textures, resulting in loss of important details in dark regions. As shown in Fig. 6, EvLight maintains more structural information but suffers from residual noise and blur artifacts. eSL-Net, with a relatively small model size, provides limited enhancement capabilities and often underperforms in challenging scenes. In contrast, our EvRWKV method achieves a better balance between noise suppression and texture preservation, producing cleaner, brighter images with rich details and minimal artifacts. This demonstrates the effectiveness of our dual-domain fusion and adaptive gating mechanisms in handling complex low-light scenarios.
In terms of quantitative evaluation, Table I lists the PSNR and SSIM scores of different methods on the SDE dataset under both SDE-in and SDE-out settings. Our method achieves the highest scores of 23.09 PSNR and 0.766 SSIM for SDE-in, as well as 23.43 PSNR and 0.767 SSIM for SDE-out. These results outperform all compared methods including SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], LLVE-SEG [35], Retinexformer [3], ELIE [17], LLVE-SEG [35], and the aforementioned qualitative comparison models. The significant improvements in both perceptual quality and objective metrics validate the superiority of our EvRWKV framework in LLIE.
@@ IV-B2 Evaluation on the SDSD Dataset
We further qualitatively compare the performance of different methods on the SDSD dataset for LLIE under challenging lighting and motion conditions. In Fig. 7, we observe that most methods improve the visibility and brightness of degraded images. However, RetinexMac [33], ELEDNet [20], and EvLight [30] exhibit limitations: As shown in Fig. 7, RetinexMac and ELEDNet tend to over-smooth textures and lose fine details, While EvLight, although preserving more structural information, causes noticeable color distortion in some regions. As shown in Fig. 7, eSL-Net [46] shows relatively limited enhancement capability due to its smaller model size. In comparison, our EvRWKV method delivers superior noise suppression and detail preservation, effectively enhancing both static and dynamic scene elements without introducing artifacts. The enhanced images show clearer textures, better contrast, and natural brightness, demonstrating the robustness of our dual-domain cross-modal fusion strategy.
@@ IV-B3 Evaluation on the RELED Dataset
To comprehensively evaluate the enhancement performance of our EvRWKV method on the RELED dataset, we compare it qualitatively with several representative approaches including eSL-Net [46], ELEDNet [20], EvLight [30], and RetinexMac [33]. As shown in Fig. 8, the RELED dataset contains extremely low-light images with severe motion blur and noise. As shown in Fig. 8, eSL-Net often causes overexposure while attempting to enhance brightness. RetinexMac boosts overall brightness but fails to eliminate noticeable blur. ELEDNet and EvLight aim to balance illumination, but this often leads to the loss of finer details. In contrast, our EvRWKV method effectively leverages event-based motion cues and image textures to suppress noise, reduce blur, and restore fine details, producing visually clearer, sharper, and more natural images without color artifacts or overexposure.
Table I reports the PSNR and SSIM scores for different methods on the RELED dataset. Our EvRWKV achieves the best quantitative results with a PSNR of 32.18 and SSIM of 0.910, significantly outperforming other methods including eSL-Net, ELEDNet, EvLight, and RetinexMac. These quantitative gains demonstrate the robustness and generalization capability of our dual-domain cross-modal fusion framework in handling challenging low-light and motion-degraded scenarios. Overall, our method achieves excellent qualitative and quantitative results across all three real-world LLIE benchmarks.
IV-C Ablation Study
We conduct comprehensive ablation studies on the SDSD-in dataset to evaluate the contribution of each key component in our EvRWKV framework. In particular, we assess the effectiveness of the EISFE module, the SpatialMix and ChannelMix mechanisms within the Cross-RWKV module, and the role of each individual loss term in the overall objective function. For each study, we remove or modify a single module while keeping the rest of the framework unchanged to isolate its impact. The results clearly demonstrate the necessity and complementarity of these components in achieving robust and high-quality LLIE. As shown in Fig. 9, the removal of EISFE (b) results in color distortion. Without SpatialMix (c), visible noise and blurred textures appear, while omitting ChannelMix (d) leads to flat tonal rendering. The full model (e) demonstrates the best visual clarity and color fidelity, especially in complex shadow regions.
IV-C1 Impact of EISFE Module
IV-C2 Impact of SpatialMix and ChannelMix
Figure 9: Visualization of ablation results.